OpenStudy (anonymous):

Medal. Solve the equation 2 over 3 (x − 2) = 4x for x. negative 2 over 5 negative 5 over 2 2 over 5 5

OpenStudy (kl0723):

\[\frac{ 2 }{ 3 }(x-2)=4x\]

OpenStudy (kl0723):

like this right?

OpenStudy (anonymous):

Yes

OpenStudy (kl0723):

2/3x-4/3=4x (2/3x-4x=4/3) whole eq times 3 -10x=4 x=4/-10 x=-2/5